# Look Around and Find Out: OOD Detection with Relative Angles

### v1

@@ 4.2 Model Ensembling with LAFO
LAFO w/ReLU 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
LAFO w/ASH 11.08 97.68 27.81 93.59 36.53 91.36 18.48 96.70 23.47 94.58
LAFO w/Scale 14.65 97.05 25.43 94.02 36.21 90.78 17.07 95.65 23.34 94.37
LAFO w/ReAct 11.13 97.79 22.34 94.95 33.33 91.81 14.65 96.60 20.36 96.29
4.3 LAFO with activation shaping algorithms
Recent methods ReAct  [38], ASH  [8] and Scale  [47] show their success to modify the feature representations to reduce model’s overconfident predictions. All three methods adopt a hyperparameter percentile to choose how to truncate and scale the feature representations using ID data statistics. When combined with Energy  [26] score, these methods improve the OOD Detection performance. On Table 3 we show that applying LAFO scoring after activation shaping algorithms improves the performance. Specifically combining LAFO with ReAct reduces FPR95 from $25.04\%$ to $20.36\%$ highlighting both the flexibility and efficacy of our method. This demonstrates that LAFO can flexibly be combined with activation shaping algorithms.
4.4 Ablation Studies
In this section, we will demonstrate the effectiveness of design choices on our score function LAFO. We first justify our choice of centering in $\bm{\mu}_{\text{ID}}$ empirically, among the candidates: $\bm{\mu}_{\text{ID}}$, $\bm{\mu}_{y_{\text{pred}}}$, $\bm{\mu}_{y_{\text{target}}}$ and $\max{(\mathbf{z}_{\text{ID}}})$. Then, we compare different angle aggregation techniques across classes by replacing our $\max{\left({\{\theta{y,y^{\prime}}\}_{y^{\prime}\in\mathcal{Y},y^{\prime}\neq y}}\right)}$ with mean and min across classes.
Table 4: Ablation on the different origin perspectives for centering. Evaluated under both ImageNet and CIFAR-10 OOD benchmarks.
@@ 4.4 Ablation Studies
Centering with $\mu_{\text{ID}}$ incorporates ID-statistics without biasing towards one particular class. Table 5 shows the performance comparison between centerings with respect to different points. Using the relative angle with respect to the predicted ($\bm{\mu}_{y_{\text{pred}}}$) or target ($\bm{\mu}_{y_{\text{target}}}$) class centroid induce a bias towards the corresponding class, which in the end hinders the compatibility between angles coming across classes. On the other hand, using $\max{(\mathbf{z}_{\text{ID}}})$ shifts every feature representation to the same orthant, reducing to simply computing the absolute distance between feature representations and the decision boundaries, which is agnostic from the in-distribution feature statistics. We observe a significant improvement in performance when computing relative angles using $\mu_{\text{ID}}$, demonstrating the importance of incorporating in-distribution (ID) statistics when measuring the relationship between feature representations and decision boundaries. LAFO with $\mu_{\text{ID}}$ centering improves the FPR95 by up to $1.45\%$ and $7.4\%$ on CIFAR-10 and ImageNet respectively while also improving the AUROC for both benchmarks.
Looking at the furthest class is better for ID/OOD separation. On Table 5 we explored different ways to aggregate class specific angles. Originally, we devise our score function to return the maximum relative angle discrepancy between the feature representation across decision boundaries. Intuitively, this suggests that considering the furthest possible class that a feature belongs from the mean of in-distribution’s perspective is effective to distinguish OOD from ID. On the other hand, comparing the minimum focuses on the smallest relative angle, reducing the separability significantly. Table 5 demonstrates taking the maximum across classes clearly outperforms mean and min aggregations, improving FPR95 and AUROC metrics on both benchmarks. Specifically the difference is higher on our large-scale experiments reducing the FPR95 by $7.72\%$ and increasing the AUROC by $1.39\%$ compared to the second best aggregation.
5 Conclusion
In this paper, we introduce a novel angle-based OOD detection score function. As a post-hoc measure of model confidence, LAFO offers several key advantages: it is (i) hyperparameter-free, (ii) model-agnostic and (iii) scale-invariant. These features allow LAFO to be applied to arbitrary pretrained models and used in conjunction with existing activation shaping algorithms, enhancing the performance. Notably, its scale-invariant nature enables simple aggregation of multiple models’ confidences through score summation, allowing a creation of an effective model ensemble for OOD detection. Our extensive experiments demonstrate that LAFO achieves state-of-the-art performance, using the relationship between the feature representations and decision boundaries relative to the ID statistics effectively. Despite the state-of-the-art performance achieved by LAFO, one possible limitation might be the use of the mean alone to capture the ID statistics in our score. As a future direction, we plan to mitigate this possible limitation by incorporating multiple relative angles to better capture the ID-statistics beyond the mean, aiming to further improve OOD detection performance. We hope that our approach inspires further research into geometric interpretations of model confidence for OOD detection.
Appendix A Appendix
A.1 Comparison of Score Distributions: CIFAR-10 vs. OOD Datasets
In this section we report the score histogram results on the Table 1 for CIFAR-10 OOD Benchmark.
A.2 Comparison of Score Distributions: ImageNet vs. OOD Datasets
We report the score histogram results on the Table 2 for ImageNet OOD Benchmark.
A.3 Model Ensemble Experiment
Table 6 show the extended version for the model ensemble experiment presented on Table 3
@@ A.3 Model Ensemble Experiment
Table 6: LAFO can be used as a score function to accumulate different architectures’ confidences due to its scale-invariance property. Evaluated under both ImageNet OOD benchmark. Best performance highlighted in bold.
@@
LAFO w/ResNet50 34.88 94.43 54.30 88.41 61.79 85.64 27.34 94.24 44.58 90.68
LAFO w/ResNet50-supcon 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
LAFO w/ViT-B/16 11.81 97.85 48.98 90.06 54.60 87.75 44.31 89.85 39.92 91.38
@@
Figure 6: Comparison of the score histograms on Imagenet (ID) and inaturalist[42](OOD) of the best individual model (left) with the model ensemble (right). Model ensemble improves the ID and OOD separation.
A.4 Implementation Details
We used Pytorch  [34] to conduct our experiments. We obtain the checkpoints of pretrained models ResNet18 with supervised contrastive loss and ResNet50 with supervised contrastive loss from  [25]’s work for a fair comparison. In the experiment where we aggregate different models’ confidences, ViT-B/16  [9] checkpoint is retrieved from the publicly available repository https://github.com/lukemelas/PyTorch-Pretrained-ViT/tree/master. In the experiment where we merge LAFO with the activation shaping algorithms ASH  [8], Scale  [47] and ReAct  [38], we used the percentiles to set the thresholds $35$, $90$ and $80$ respectively. All experiments are evaluated on a single Nvidia H100 GPU. Note that, due to our hyperparameter-free post-hoc score function, all experiments are deterministic given the pretrained model.

### v2

@@ A.2 Comparison with ReAct
ReAct 20.38 96.22 24.20 94.20 33.85 91.58 47.30 89.80 31.43 92.95
ORA 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
ORA w/ReAct 11.13 97.79 22.34 94.95 33.33 91.81 14.65 96.60 20.36 96.29
Table 11: ORA vs ReAct under CIFAR OOD benchmark.
Method SVHN iSUN Places Texture Avg
@@
ReAct 6.15 98.75 10.31 98.09 21.68 95.47 10.18 98.12 12.08 97.61
ORA 3.53 99.16 8.36 98.28 23.40 94.88 8.58 98.34 10.97 97.67
ORA w/ReAct 3.35 99.18 8.11 98.29 20.84 95.25 7.87 98.45 10.04 97.79
A.3 Centering with Different Statistics
Tables 12 and 13 present the performance of the ORA score when using different class means as the reference view instead of the mean of ID features. We also explored alternative centering strategies by replacing the mean of ID features with elementwise operations—max, min, and median—where each corresponds to using the respective statistic before calculating angles. Additionally, sum aggregation refers to summing scores obtained from individual class mean reference points or from these elementwise operations. Results indicate that using $\mu_{\text{ID}}$ consistently outperforms these alternatives.
Table 12: CIFAR10 centering with different statistics using ResNet18 model.
Method SVHN iSUN Places Texture Avg ID Acc
@@ A.3 Centering with Different Statistics
Class 0 mean 4.77 98.96 8.06 98.27 25.20 94.62 10.11 98.19 12.03 97.51 95.2
Class 1 mean 6.12 98.77 8.86 98.24 24.94 94.96 13.16 97.68 13.27 97.41 98.5
Class 2 mean 5.42 98.84 7.90 98.36 22.19 95.58 11.42 97.98 11.73 97.69 92.6
Class 3 mean 5.94 98.76 7.99 98.29 22.80 95.43 11.35 97.66 12.02 97.54 88.7
Class 4 mean 5.44 98.85 8.87 98.22 22.68 95.47 11.26 97.96 12.06 97.63 95.8
Class 5 mean 6.22 98.64 7.38 98.45 23.11 95.48 11.84 97.82 12.14 97.60 88.0
Class 6 mean 5.74 98.82 8.50 98.26 97.67 20.76 12.11 95.73 11.78 97.62 97.9
Class 7 mean 5.93 98.78 8.29 98.30 24.55 95.13 12.57 97.82 12.84 97.51 96.4
Class 8 mean 5.81 98.81 10.03 97.95 26.79 94.18 10.41 98.11 13.26 97.26 97.0
Class 9 mean 6.11 98.78 9.00 98.19 24.89 94.62 11.35 97.95 12.84 97.25 96.3
Sum aggregation 5.68 98.85 8.27 98.33 23.70 95.34 11.33 97.98 12.25 97.62 -
Elementwise max 6.28 98.73 8.28 98.30 13.79 97.57 24.35 95.21 13.18 97.45 -
Elementwise min 3.60 99.14 14.82 97.10 9.38 97.99 27.62 92.97 13.85 96.80 -
Elementwise median 2.33 99.34 10.02 97.90 7.73 98.29 23.99 93.84 11.02 97.34 -
Sum aggregation 5.78 98.65 20.31 95.64 10.35 97.80 30.42 91.60 16.72 95.92 -
ORA 3.53 99.16 8.36 98.28 23.40 94.88 8.58 98.34 10.97 97.67 94.6
Table 13: ImageNet centering with different statistics using ResNet50 model.
Method iNaturalist SUN Places Texture Avg ID Acc
[…]
Class 1 mean 16.01 96.92 31.63 92.52 39.86 90.67 25.39 93.33 28.22 93.36 92
Class 250 mean 11.48 97.65 31.20 92.69 39.53 90.77 20.16 94.87 25.59 93.99 66
Class 500 mean 14.87 97.08 38.97 90.52 45.80 88.90 26.29 93.04 31.48 92.28 60
Class 750 mean 11.57 97.59 34.23 92.13 42.60 90.15 19.75 95.18 27.04 93.76 84
Class 1000 mean 11.36 97.63 30.20 93.01 38.12 91.08 19.31 95.31 24.75 94.26 60
Sum aggregation 12.40 97.48 32.47 92.36 40.42 90.53 21.38 94.52 26.67 93.72 -
Elementwise Max 17.10 96.76 34.22 91.73 41.88 90.14 35.09 89.84 32.07 92.12 -
Elementwise Min 29.16 94.51 60.70 85.81 65.01 83.18 22.84 95.07 44.43 89.64 -
Elementwise Median 20.04 95.83 46.66 89.15 54.42 85.61 15.04 96.81 34.04 91.85 -
Sum aggregation 21.09 95.89 49.47 88.97 56.50 86.12 17.62 95.98 36.17 91.74 -
ORA 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26 77.33
A.4 Resource Constrained Setting
Table 14 presents the performance of ORA in a resource-constrained setting using MobileNetV2 Sandler et al. [2018], a model designed for efficient inference with minimal resources. ORA achieves the best average performance, improving FPR and AUROC by $7.56\%$ and $1.66\%$, respectively, demonstrating its adaptability to resource-limited scenarios.
Table 14: Resource Constrained Setting: ImageNet MobileNet_v2 performances.
Method iNaturalist SUN Places Texture Avg
[…]
MSP 59.84 86.71 74.15 78.87 76.84 78.14 70.98 78.95 70.45 80.67
Energy 55.35 90.33 59.36 86.24 66.28 83.21 54.54 86.58 58.88 86.59
KNN 85.92 72.67 90.51 65.39 93.21 60.08 14.04 96.98 70.92 73.78
fDBD 53.72 90.89 68.22 82.84 73.20 80.09 37.82 91.85 58.24 86.42
ORA 46.59 91.86 61.21 85.01 67.81 82.08 27.07 94.04 50.68 88.25
Appendix B Implementation Details
We used Pytorch  [Paszke et al., 2019] to conduct our experiments. We obtain the checkpoints of pretrained models ResNet18 with supervised contrastive loss and ResNet50 with supervised contrastive loss from  Liu and Qin [2024]’s work for a fair comparison. In the experiment where we aggregate different models’ confidences, ViT-B/16  [Dosovitskiy et al., 2020] checkpoint is retrieved from the publicly available repository https://github.com/lukemelas/PyTorch-Pretrained-ViT/tree/master. In the experiment where we merge ORA with the activation shaping algorithms ASH  [Djurisic et al., 2023], Scale  [Xu et al., 2024b] and ReAct  [Sun et al., 2021], we used the percentiles to set the thresholds $35$, $90$ and $80$ respectively. For the extended results on Table 1, we used the timm [Wightman, 2020] checkpoints for the models ConvNeXt [Liu et al., 2022], Swin [Liu et al., 2021], DeiT [Touvron et al., 2021] and EVA [Fang et al., 2023]. Similarly, for the CLIP experiments on Table 3, we used the huggingface checkpoint of CLIP ViT-H/14 [LAION, ]. All experiments are evaluated on a single Nvidia H100 GPU. Note that, thanks to our hyperparameter-free post-hoc score function, all experiments are deterministic given the pretrained model.
Appendix C Robustness of the Relative Angles
In this section, we provide a theoretical justification for focusing on relative angles, rather than absolute distances, to better understand the robustness of representations under scaling transformations.
Theorem: Let $M_{1}$ and $M_{2}$ be two neural networks such that:
  1. 1.
    The encoder of $M_{2}$ is a scaled version of the encoder of $M_{1}$. Specifically, there exists a positive scalar $k\in\mathbb{R}_{+}$ such that the output of the penultimate layer of $M_{2}$ satisfies:
    $\mathbf{z}^{M_{2}}=k\cdot\mathbf{z}^{M_{1}}$
  2. 2.
    Both networks share the same final linear layer (without bias) for classification.
Since scaling transformations do not affect the softmax decision boundaries (due to monotonicity), $M_{1}$ and $M_{2}$ will produce the same classification decisions. Under this setup:
  • The angles satisfy $\theta_{y_{1},y_{2}}(\mathbf{z}^{M_{1}})=\theta_{y_{1},y_{2}}(\mathbf{z}^{M_{2}})$, where $\theta_{y_{1},y_{2}}(\mathbf{z})$ is the relative angle between representation and its projection onto the decision boundary.
  • However, the distances satisfy $d_{y_{1},y_{2}}(\mathbf{z}^{M_{2}},\mathbf{z}_{db}^{M_{2}})=k\cdot d_{y_{1},y_{2}}(\mathbf{z}^{M_{1}},\mathbf{z}_{db}^{M_{1}})$, where $d(\cdot,\cdot)$ is the Euclidean distance.
Proof: Let $\mathbf{z}^{M_{1}}$ be the penultimate layer representation of $M_{1}$ and let $\mathbf{z}_{db}^{M_{1}}$ denote the projection of $\mathbf{z}^{M_{1}}$ onto the decision boundary between two classes $y_{1}$ and $y_{2}$. Then following the Equation 1
$\mathbf{z}_{db}^{M_{1}}=\mathbf{z}^{M_{1}}-\frac{(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})^{T}\mathbf{z}^{M_{1}}}{\|\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}}\|^{2}}(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})$
where $\mathbf{w}_{y_{1}}$ and $\mathbf{w}_{y_{2}}$ are the weight vectors corresponding to classes $y_{1}$ and $y_{2}$ in the shared linear layer. For $M_{2}$, the features are scaled by $k$, so $\mathbf{z}^{M_{2}}=k\cdot\mathbf{z}^{M_{1}}$. Substituting into the projection formula:
$\mathbf{z}_{db}^{M_{2}}$ $=k\cdot\mathbf{z}^{M_{1}}-\frac{(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})^{T}k\cdot\mathbf{z}^{M_{1}}}{\|\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}}\|^{2}}(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})$
$\mathbf{z}_{db}^{M_{2}}$ $=k\cdot\mathbf{z}_{db}^{M_{1}}$
The angle is defined by the cosine similarity between the vectors $\mathbf{z}^{M_{1}}-\mathbf{\mu}_{\text{ID}}^{M_{1}}$ and $\mathbf{z}_{db}^{M_{1}}-\mathbf{\mu}_{\text{ID}}^{M_{1}}$. Using the substitutions $\mathbf{z}^{M_{2}}=k\cdot\mathbf{z}^{M_{1}}$, $\mathbf{z}_{db}^{M_{2}}=k\cdot\mathbf{z}_{db}^{M_{1}}$, and $\bm{\mu}_{\text{ID}}^{M_{2}}=k\cdot\bm{\mu}_{\text{ID}}^{M_{1}}$, we have:
$\cos{(\theta_{y_{1},y_{2}}(\mathbf{z}^{M_{1}}))}$ $=\frac{\langle\mathbf{z}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}},\mathbf{z}_{db}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\rangle}{\|\mathbf{z}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\|\cdot\|\mathbf{z}_{db}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\|}$
$=\frac{k\cdot k\cdot\langle\mathbf{z}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}},\mathbf{z}_{db}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\rangle}{k\cdot k\cdot\|\mathbf{z}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\|\cdot\|\mathbf{z}_{db}^{M_{1}}-\bm{\mu}_{\text{ID}}^{M_{1}}\|}$
$=\frac{\langle k\cdot\mathbf{z}^{M_{1}}-k\cdot\bm{\mu}_{\text{ID}}^{M_{1}},k\cdot\mathbf{z}_{db}^{M_{1}}-k\cdot\bm{\mu}_{\text{ID}}^{M_{1}}\rangle}{\|k\cdot\mathbf{z}^{M_{1}}-k\cdot\bm{\mu}_{\text{ID}}^{M_{1}}\|\cdot\|k\cdot\mathbf{z}_{db}^{M_{1}}-k\cdot\bm{\mu}_{\text{ID}}^{M_{1}}\|}$
$=\frac{\langle\mathbf{z}^{M_{2}}-\bm{\mu}_{\text{ID}}^{M_{2}},\mathbf{z}_{db}^{M_{2}}-\bm{\mu}_{\text{ID}}^{M_{2}}\rangle}{\|\mathbf{z}^{M_{2}}-\bm{\mu}_{\text{ID}}^{M_{2}}\|\cdot\|\mathbf{z}_{db}^{M_{2}}-\bm{\mu}_{\text{ID}}^{M_{2}}\|}$
$=\cos{(\theta_{y_{1},y_{2}}(\mathbf{z}^{M_{2}}))}$
On the other hand, the absolute distance between $z$ and $z_{db}$ scale as follows:
$d_{y_{1},y_{2}}(\mathbf{z}^{M_{2}},\mathbf{z}_{db}^{M_{2}})$ $=d_{y_{1},y_{2}}(k\cdot\mathbf{z}^{M_{1}},k\cdot\mathbf{z}_{db}^{M_{1}})$
$=k\cdot d_{y_{1},y_{2}}(\mathbf{z}^{M_{1}},\mathbf{z}_{db}^{M_{1}})$
Therefore, for two networks $M_{1}$ and $M_{2}$ with identical performance, we demonstrate that relative angles remain invariant to scaling, whereas absolute distances are sensitive to it. Given that ReLU networks are commonly used in practice (where activations are unbounded), scale-invariant, angle-based techniques provide a more robust and suitable approach for measuring confidence compared to distance-based methods, especially when comparing the confidences of different models.
Appendix D Algorithm Box
We present the pseudocode for ORA in Algorithm Box 1. It depicts how ORA assigns a score given a sample $x$, pretrained model $f$, and the ID statistics, mean of the in-distribution features $\mu_{\text{ID}}$.
Algorithm 1 ORA (OOD Detection with Relative Angles)
2:OOD score $s$
3:function ORA($\mathbf{x},f,\bm{\mu}_{\text{ID}}$)
5:  $\mathbf{z}=f_{1}\circ\ldots\circ f_{L-1}(\mathbf{x})$ $\triangleright$ penultimate layer features
6:  $score\leftarrow$-$\infty$
8:   compute $\mathbf{z}_{db}$ as in Eq.  1
9:   compute $\theta_{\hat{y},y^{\prime}}(\mathbf{z})$ using Eq.  2
10:   compute $\tilde{s}(\mathbf{z})$ using Eq.  3
11:   if $\tilde{s}(\mathbf{z})\geq score$ then
12:     $score=\tilde{s}(\mathbf{z})$
13:   end if
14:  end for
15:  return scorescore $\triangleright$ maximum score across classes
16:end function
Appendix E Detailed Tables and Figures
@@
Table 15: Extended version for the model ensemble experiment presented on Table 4. ORA can be used as a score function to accumulate different architectures’ confidences due to its scale-invariance property. Evaluated under both ImageNet OOD benchmark. Best performance highlighted in bold.
@@
ORA w/ResNet50 34.88 94.43 54.30 88.41 61.79 85.64 27.34 94.24 44.58 90.68
ORA w/ResNet50-supcon 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
ORA w/ViT-B/16 11.81 97.85 48.98 90.06 54.60 87.75 44.31 89.85 39.92 91.38
@@
Figure 3: Comparison of the score histograms on Imagenet (ID) and inaturalist[Van Horn et al., 2018](OOD) of the best individual model (left) with the model ensemble (right). Model ensemble improves the ID and OOD separation.
Table 16: Extended version for the activation shaping experiment presented on Table 5. ORA can be used as a plug-in on top of activation shaping algorithms. Evaluated under ImageNet OOD benchmark. $\uparrow$ indicates that larger values are better and vice versa. Best performance highlighted in bold.
Method iNaturalist SUN Places Texture Avg
[…]
ORA w/ReLU 12.27 97.42 31.80 92.85 40.71 90.10 15.39 96.68 25.04 94.26
ORA w/ASH 11.08 97.68 27.81 93.59 36.53 91.36 18.48 96.70 23.47 94.58
ORA w/Scale 14.65 97.05 25.43 94.02 36.21 90.78 17.07 95.65 23.34 94.37
ORA w/ReAct 11.13 97.79 22.34 94.95 33.33 91.81 14.65 96.60 20.36 96.29
Appendix F Plain Model Performances
Tables 17, 18, and 19 show the performance of feature-based OOD methods on models trained without supervised contrastive loss. The results highlight that supervised contrastive loss significantly enhances feature quality, leading to a substantial performance boost for feature-based OOD methods.
Table 17: CIFAR10 Plain ResNet18 performances.
Method SVHN iSUN Places Texture Avg
[…]
KNN 27.85 95.52 24.67 95.52 44.56 90.85 37.57 94.71 33.66 94.15
fDBD 22.58 96.07 23.96 95.85 46.59 90.40 31.24 94.48 31.09 94.20
ORA 22.09 96.02 22.91 95.90 46.46 90.37 31.28 94.48 30.86 94.21
Table 18: ImageNet Plain ResNet50 performances.
Method iNaturalist SUN Places Texture Avg
[…]
KNN 59.00 86.47 68.82 80.72 76.28 75.76 11.77 97.07 53.97 85.01
fDBD 40.24 93.67 60.60 86.97 66.40 84.27 37.50 92.12 51.19 89.26
ORA 38.94 93.68 59.78 86.53 66.89 83.04 31.67 93.33 49.32 89.15
Table 19: ImageNet ViT performances.
Method iNaturalist SUN Places Texture Avg
[…]
KNN 11.41 97.65 56.91 86.39 63.76 82.61 42.23 89.61 43.58 89.07
fDBD 12.86 97.72 50.86 89.74 56.28 87.44 45.74 89.41 41.44 91.08
ORA 11.80 97.86 48.81 90.14 54.32 87.88 44.56 89.75 39.87 91.41
Appendix G Histogram Plots for ID/OOD Separability
Figures 4 and 5 shows the score distributions on the Tables 8 and 9 respectively.
Appendix H Comparison of Different Angles
Figure 6 presents the score distributions for different angles discussed in Section 3.1. The results show that $\sin{(\alpha)}$ does not provide good ID/OOD separation, whereas $\sin{(\theta)}$ and $\sin{(\theta)}/\sin{(\alpha)}$ present significantly clearer distinctions. Additionally, incorporating $\sin{(\alpha)}$ into $\sin{(\theta)}/\sin{(\alpha)}$ hinders its performance compared to using $\sin{(\theta)}$ alone.
Figure 6: We demonstrate the ID/OOD separability of $\sin(\alpha)$, $\sin(\theta)$ and $\frac{\sin(\theta)}{\sin(\alpha)}$. Columns show the performances on iNaturalist and SUN datasets respectively. It can be seen that the ID/OOD class separability is the best when $\sin(\theta)$ is used: considering $\sin(\alpha)$ impedes the performance as confirmed quantitatively in terms of FPR95 and AUROC metrics in Table 9.
Appendix I Near OOD Results
In Table 20, we report ORA’s FPR and AUROC on NINCO [Bitterwolf et al., 2023] and SSB-hard [Vaze et al., 2022] datasets, along with average scores and in-distribution (ID) classification accuracy.
NINCO dataset: We observe that models pretrained on ImageNet21k generally outperform their non-pretrained counterparts in terms of FPR95, consistent with the observations in Bitterwolf et al. [2023]. However, no single method, including ORA, dominates across all architectures. For example, ORA attains the best results on the plain ConvNeXt Liu et al. [2022] backbone, while Rcos [Bitterwolf et al., 2023] outperforms others on DeiT [Touvron et al., 2021] and ResNet50 [He et al., 2016]. Interestingly, Energy [Liu et al., 2020] performs best on Swin-pretrained [Liu et al., 2021].
SSB-Hard dataset: A similar trend is evident: ORA outperforms other methods on plain ConvNeXt [Liu et al., 2022] and offers strong performance on SWIN-pre [Liu et al., 2021], ConvNeXt-pre [Liu et al., 2022], and DeiT-pre [Touvron et al., 2021], where it outperforms Rcos [Bitterwolf et al., 2023], RMaha [Ren et al., 2021], and Pcos [Galil et al., ]. However, it is worse than naive baselines like Energy [Liu et al., 2020] in certain pretrained settings. In contrast, for models without pre-training, naive methods degrade significantly while ORA remains on par with the top baselines.
Discussion. These findings underscore a fundamental distinction between far- and near-OOD detection. Identifying completely novel inputs (as in far-OOD) is a fundamentally different problem from determining how much deviation is acceptable within or between known classes (as in near-OOD). Consequently, methods designed for detecting novelty or semantic outliers—such as those based on global uncertainty or energy scores—do not necessarily excel in near-OOD settings. This makes direct comparisons across these tasks potentially misleading. In near-OOD scenarios, where fine-grained class separation is critical, approaches that explicitly incorporate class-wise feature statistics—such as Rcos, Pcos, or RMah [Bitterwolf et al., 2023, Galil et al., , Ren et al., 2021]—tend to have a clear advantage. These methods are better equipped to handle intra-class variance and subtle distributional shifts, highlighting the need for tailored evaluation and method design across OOD subtypes.
[…]
Model NINCO SSB-Hard Average ID Acc (%)
[…]
ResNet-50 68.31 85.89 85.65 71.17 76.98 78.53 76.12
ResNet-50-SCL 54.39 86.50 79.09 71.35 66.74 78.93 77.31
ConvNeXt 48.04 89.17 70.90 75.47 59.47 82.32 84.84
ConvNeXt-pre 45.54 90.56 68.82 77.13 57.18 83.84 85.49
DeiT 71.76 78.47 83.07 69.38 77.42 73.93 83.13
DeiT-pre 60.07 86.40 80.40 72.40 70.24 79.40 84.79
Swin 66.91 84.00 81.28 73.95 74.10 78.98 84.49
Swin-pre 52.62 88.25 77.45 74.76 65.04 81.51 85.73
EVA 46.61 88.15 67.86 75.65 57.24 81.90 87.88
Appendix J Relation to [Liu and Qin, 2024]
In this section we are going to describe the methodology relation of our method with respect to the method of Liu and Qin [2024], explaining the performance gap in favor of ORA.
Let $\mathbf{z}\in\mathbb{R}^{D}$ be a feature, $\mathbf{z}_{db}$ its orthogonal projection onto the decision boundary given in (1), and let $\bm{\mu}_{\text{ID}}\in\mathbb{R}^{D}$ denote the in-distribution mean. Define centered quantities
$\tilde{\mathbf{z}}=\mathbf{z}-\bm{\mu}_{\text{ID}},\qquad\tilde{\mathbf{z}}_{db}=\mathbf{z}_{db}-\bm{\mu}_{\text{ID}}.$
In the triangle with vertices $\{\mathbf{0},\tilde{\mathbf{z}},\tilde{\mathbf{z}}_{db}\}$, the law of sines gives
$\|\mathbf{z}-\mathbf{z}_{db}\|\;=\;\|\tilde{\mathbf{z}}\|\;\frac{\sin\!\big(\theta\big)}{\sin\!\big(\alpha\big)},$ (5)
where $\theta$ is the angle at $\tilde{\mathbf{z}}$ (discriminative ID/OOD signal, measured in the ID-centered frame) and $\alpha$ is the auxiliary angle at the origin $\mathbf{0}$; the radial term $\|\tilde{\mathbf{z}}\|$ captures nuisance scale (e.g., global brightness/contrast).
The fDBD score normalizes out the scale term:
$\mathrm{fDBD}(\mathbf{z})=\frac{\|\mathbf{z}-\mathbf{z}_{db}\|}{\|\tilde{\mathbf{z}}\|}=\frac{\sin\!\big(\theta\big)}{\sin\!\big(\alpha\big)},$
which makes the score invariant to uniform rescaling of features and improves over raw distance.
Empirically, $\sin\!\big(\alpha\big)$ is nearly identical for ID and OOD (see Fig. 2), adding variance without separation. ORA therefore retains only the informative angular component in the ID-centered frame:
$\mathrm{s}_{\text{ORA}}(\mathbf{z})\;\propto\;\theta$
Hierarchy.
Raw distance $d=\|\mathbf{z}-\mathbf{z}_{db}\|$ is confounded by the ID-centered norm $\|\tilde{\mathbf{z}}\|=\|\mathbf{z}-\bm{\mu}_{\text{ID}}\|$. The fDBD score removes this scale confound but still carries the factor $1/\sin\!\alpha$. ORA discards that factor and retains only the discriminative angle, which explains its per-model gains and its suitability for scale-invariant ensembling.
Appendix K Results on small datasets
To confirm empirically the hypothesis that ID points lie further from the decision boundary holds even for simple datasets and models, we trained a LeNet-5 [LeCun et al., 2002] architecture on MNIST and measured the distance to the decision boundary for MNIST (ID) and Fashion-MNIST (OOD).
Dataset Mean Std Min Max
MNIST 5.86 0.62 2.69 7.49
Fashion-MNIST 4.77 0.82 1.99 6.66
Table 21: Distance to the decision boundary for MNIST (ID) and Fashion-MNIST (OOD).
ID samples are, on average, 1.1 units further from the decision boundary than OOD samples.
Discussion.
Cross-entropy training maximizes class-conditional log-likelihood, which drives features to align strongly with their corresponding weight vectors (confident regions). This places ID features deep inside their class cones and therefore further from the boundary. In contrast, OOD images do not match strong class-specific patterns, so their features align less and end up closer to the boundary, as reflected in the distance statistics.
This experiment supports empirically the assumption on the distribution of in-distribution points with respect to the decision boundaries which is at the core of our method and fDBD Liu and Qin [2024].
Appendix L Future Work
In this paper, we explored several alternative reference frames for the angle computation, including centering with respect to each individual class mean and centering only on the predicted class’s mean. However, neither variant yielded consistent improvements over our global-mean formulation. A natural next step is to move beyond this either–or choice toward a weighted combination of per-class angle scores. Such weights could be derived from the ID data. For instance, inversely proportional to the in-distribution feature variance of each class or proportional to the class-wise ID accuracy. This approach would preserve ORA’s hyperparameter-free character when weights are computed from readily available ID statistics, while allowing the detector to adaptively emphasize the most reliable class anchors. We leave a systematic study of this direction for future work.